# On Leakage of Code Generation Evaluation Datasets

### v2

@@ 4 Conclusion
We study the cause and effect of data contamination by two popular code generation benchmarks. Our analysis highlights that contamination is likely unavoidable at the LLM scale given the difficulty of filtering every potential permutation of a benchmark dataset. This insight motivates our contribution of a novel code generation benchmark to evaluate contemporary LLMs in a contamination-free setting. We are well aware that our decision to release this dataset will make future leakage impossible to control. However, with the context of the fast-paced model development cycles that LLMs are currently undergoing we believe that releasing this increases trustworthiness and usefulness of this dataset. It is conveniently designed to serve as drop-in replacement (or addition) of current evaluation sets. On top of its newness, the more challenging nature of this dataset also allows it to provide more signal for model comparison.
@@ 5 Limitations
All the model analysis was done black-box, without inspecting the model weights or the training set (excepting the work on synthetic data). There is no reason why this dataset will not follow the same path than the two studied here. As mentioned in the Conclusion we believe there is more value in that than in an alternative solution (not releasing or keeping it behind an API access).
Appendix
@@
Write a function to find the perimeter of a square Calculate the perimeter of a square
Write a python function to find the volume of a triangular prism How can I use JavaScript to calculate the volume of a triangular prism with an irregular base and uneven side lengths?
Write a function to convert snake case string to camel case string. Convert the sentence to camel case.
Table 1: Examples of similar prompts from the test set of MBPP found in evol_instruct (Xu et al., 2023).
@@
Table 3: Most similar prompt in evol-instruct for a non-handpicked selection of HumanEval prompts.
@@
Model Name Family pass@1
Mistral-7b Mistral 0.11
Mixtral 8x7B Mistral 0.17
Mixtral 8x22B Mistral 0.38
Codestral 22B Mistral 0.40
Codellama-7b-instruct Meta 0.14
Codellama-34b-instruct Meta 0.19
Llama2-7b-chat Meta 0.02
Llama2-70b-chat Meta 0.10
Llama3-8b-instruct Meta 0.27
Llama3-70b-instruct Meta 0.49
GPT-3.5-tb OpenAI 0.37
GPT-4 OpenAI 0.53
GPT-4o OpenAI 0.63
Claude-2 Anthropic 0.34
Claude-3-haiku Anthropic 0.34
Claude-3-sonnet Anthropic 0.40
Claude-3-opus Anthropic 0.54
Deepseek-coder-33b-instruct DeepSeek 0.39
DBRX Instruct DBRX 0.25
Qwen1.5-72B-Chat Qwen 0.20
Qwen1.5-110B-Chat Qwen 0.30
Qwen 2-72B Instruct Qwen 0.42
Command R+ Cohere 0.22
Command R Cohere 0.12
Table 4: Results of pass@1 rate on the LBPP dataset for a selection of models.

### v3

@@ 5 Conclusion
We study the cause and effect of data contamination via two popular code generation benchmarks. Our analysis highlights that contamination is likely unavoidable at the LLM scale given the difficulty of filtering every potential permutation of a benchmark dataset. This insight motivates our contribution of LBPP: a novel code generation benchmark to evaluate contemporary LLMs in a contamination-free setting. We are well aware that our decision to release this dataset will make future leakage impossible to control. However, with the context of the fast-paced model development cycles that LLMs are currently undergoing we believe that releasing this increases the trustworthiness and usefulness of this dataset. It is conveniently designed to serve as a drop-in replacement (or addition) of current evaluation sets. On top of its novelty, the more challenging nature of this dataset also provides a cleaner signal for model comparison.
@@ 6 Limitations
All the model analysis was done black-box, without inspecting the model weights or the training set (except the work on synthetic data). There is no reason why this dataset will not follow the same path as the two studied here. As mentioned in the Conclusion we believe there is more value in that than in an alternative solution (not releasing or keeping it behind an API access).
Appendix A Appendix
@@
Write a function to find the
perimeter of a square
Calculate the perimeter of a square
Write a python function to find
the volume of a triangular prism
How can I use JavaScript to
calculate the volume of a
triangular prism with an irregular
base and uneven side lengths?
Write a function to convert snake
case string to camel case string.
Convert the sentence to camel case.
Table 3: Similarity in prompts between MBPP evaluation dataset and evol-instruct synthetic training dataset.
@@
Table 5: Most similar prompt in evol-instruct for a random sample of HumanEval prompts.